# Vortices and Factorization

### v2

@@ 11 Equilibria of Vortex Streets - Even Family
In the previous section we considered vortex street configurations related to the trigonometric analogs of the odd family of bi-spectral potentials or soliton solutions of the KdV hierarchy. Here we consider trigonometric analogs of the even family (see section 6), i.e. examples of periodic configurations where not all ${\cal Q}_{i}$ are integers. For this, let us start our chain of Darboux transformations with the trigonometric Pochl-Teller potential
$u_{0}=\frac{a(a-1)}{\sin^{2}z}+\frac{b(b-1)}{\cos^{2}z}$ (83)
with arbitrary real $a$ and $b$. The simplest configuration corresponds to the eigenfunction $\sin^{a}z\cos^{b}z$ and consists of two static vortices of the strengths ${\cal Q}_{1}=a$, ${\cal Q}_{2}=b$ situated at the diametrically opposite points of the $\pi$-periodic cylinder, i.e. at $z_{1}=0$, $z_{2}=\pi/2$.
Note that when $a=b$, or $a=1-b$, the potential (83) can be rewritten as $u_{0}=4a(a-1)/\sin^{2}2z$. By re-scaling $2z\to z$ in the Schrodinger operator, we again come to the special $b=0$ case of the $\pi$-periodic potential (83).
For integers $a$ and $b$, (83) is a soliton solution of KdV at specific “times” (corresponding parameters $k_{i}$ and $\zeta_{i}$ in (79, 78) can be found e.g. in refs [6] or [18]) [^13] In general, sequence of $n$ Darboux transformations of the chain (34), where $A_{i}=\partial_{z}-(a+i)\cot z+(b+i)\tan z$, $i=0,1,\dots$., results in the shift of parameters $a\to a+n$, $b\to b+n$ in the potential (83).. In this case, $u_{0}$ and sequences of its Darboux transforms belong to the odd family considered in the previous section.
Due to invariance of the potential (83) wrt involutions $a\to 1-a$, $b\to 1-b$, when $a\pm b$ is not integer (generic $a$ and $b$), the Schrodinger operator $H_{0}=-\partial_{z}^{2}+u_{0}$ has four distinct types of factorizable eigenfunctions $\varkappa_{i}(z)$, $H_{0}\varkappa_{i}=\lambda_{i}\varkappa_{i}$ of the form
@@
where $i=0,1,2,\dots$ and ${}_{2}F_{1}$ stands for the Gauss hypergeometric function. The first parameter of the ${}_{2}F_{1}$ in (84) is a non-positive integer $-i$, and the hypergeometric factors in (84) are polynomials in $\sin^{2}z$. These factorizable eigenfunctions are trigonometric-rational up to a factor $\sin^{\pm a}z\cos^{\pm b}z$. For generic $a$ and $b$ all eigenvalues $\lambda_{i}$ are distinct and all eigenfunctions in (84) are linearly independent.
From (84), with the help of the Wronskian formula (79), we can construct eigenfunctions of Darboux transforms $H_{n}=-\partial_{z}^{2}+u_{0}-2(\log\tau_{n})^{\prime\prime}$ of $H_{0}$. These eigenfunctions are factorizable, and therefore correspond to static vortex street configurations. Note that, for generic $a$ and $b$, only one of two linearly independent solutions of the Schrodinger equation $H_{0}\varkappa_{i}=\lambda_{i}\varkappa_{i}$ is factorizable. Therefore, in difference from the soliton-related configurations (79,78), configurations generated by (84) do not depend on free parameters others than $a$ and $b$. Similarly to systems considered in section 9, configurations with generic $a$ and $b$ are parametrized only by sequences of distinct integers $0<k_{1}<k_{2}<\dots<k_{n}$.
Even family of periodic solutions emerges when $a$ and $b$ become half-integers, such that
@@
In this case, eigenvalues $\lambda_{i}$ in (84) overlap, and some of them become degenerate, with two-dimensional eigenspaces [^14] Degeneration occurs also for integer $a$ and $b$. There, number of degenerate states, as we saw in the previous section, is infinite.. Now, there are two types (degenerate and non-degenerate) of eigenfunctions $\varkappa_{i}(z)$ :
$\begin{array}[]{lll}P^{(-l,-m)}_{i}(\cos 2z,\zeta_{i})\sin^{1/2-l}z\cos^{1/2-m}z,&\lambda_{i}=(2i+1-l-m)^{2},&\frac{l+m}{2}\leq i<l+m\\ P^{(l,m)}_{i}(\cos 2z)\sin^{l+1/2}z\cos^{m+1/2}z,&\lambda_{i}=(2i+1+l+m)^{2},&i\geq 0\end{array},$ (85)
which are both trigonometric-rational up to the common factor $(\sin z\cos z)^{1/2}$. The first type, $P^{(-l,-m)}_{i}(x,\zeta)$ is a polynomial in $x$ which depending also on a free parameter $\zeta$. Such polynomials are called para-Jacobi Polynomials [11], [8]. Solutions of the second type involve ordinary Jacobi polynomials $P^{(l,m)}_{i}$ and are non-degenerate. In difference from the latter, the para-Jacobi polynomials are not orthogonal.
@@
The simplest non-trivial example of a member of the even family corresponds to $l=m=1$. Here $a=b=3/2$, $P^{(-1,-1)}_{1}(\cos 2z,\zeta_{1})=\cos 2z+1+\zeta_{1}$ and from (85) we get
@@
Note that by the scaling $z\to\epsilon z$, $\psi_{1}\to-\psi_{1}/(2\epsilon^{3/2})$ and change of parameter $\zeta_{1}=-2(1+\epsilon^{2}t_{1})$, in the $\epsilon\to 0$ limit, we obtain $\psi_{1}=(z^{2}+t_{1})/z^{1/2}$. This is $\psi_{1}$ (see (56)) of the even bi-spectral family of section 6.
@@
$\varkappa_{m+i}=(\sin z\cos z)^{1/2-m}P^{(-m,-m)}_{m+i}(\cos 2z,\zeta_{m+i}),\quad i=0,1,\dots,m-1$ (87)
@@
When we re-scale $z\to\epsilon z$, the eigenvalues scale as $\lambda_{i}\to\epsilon^{2}\lambda_{i}$, and in the rational limit we have confluent eigenvalues, all going to $\lambda_{i}\to\eta=0$. According to section 4 (see eqs. (42), (43)) in this limit the Wronskian representation of the Darboux transforms becomes
@@
Here $H_{0}=-\partial_{z}^{2}+\frac{m^{2}-1/4}{z^{2}}$ is the rational limit of the trigonometric Schrodinger operator (83) with $a=b=m+1/2$, obtained by the scaling $z\to\epsilon z$, $H_{0}\to\epsilon^{-2}H_{0}$. In other words, ${\cal X}_{i}(z)$ are Laurent polynomials in $z^{1/2}$ that solve the chain
${\cal X}_{1}=z^{1/2-m}(z^{2m}+t_{m}),\quad-{\cal X}_{i+1}^{\prime\prime}+\frac{m^{2}-1/4}{z^{2}}{\cal X}_{i+1}=c_{i}{\cal X}_{i},\quad i=1,2,\dots,m-1,$ (89)
where $c_{i}$ are arbitrary constants. Although two constants of integration appear at each step of the chain, only one of them is essential. The chain terminates at the $(m-1)$th step, since the logarithmic term appears at the $m$th step. Function ${\cal X}_{i}$ is the $\epsilon\to 0$ limit of a linear combinations of $i$ first eigenfunctions (87):
@@
$1,\quad\epsilon^{2m}t_{m},\quad\epsilon^{2m+2}t_{m+1},\quad\epsilon^{2m+4}t_{m+1},\dots\epsilon^{2(m+j-1)}t_{m+j}.$
Coefficients of linear combinations are constants determined by the condition of existence of the limit.
For example, in $m=2$ case, corresponding to the terminating sequence (57), we have
$\begin{array}[]{l}{\cal X}_{1}=\lim_{\epsilon\to 0}\epsilon^{-5/2}\varkappa_{2}(\epsilon z,2-2\epsilon^{4}t_{2}),\\ {\cal X}_{2}=\lim_{\epsilon\to 0}\epsilon^{-9/2}\left(\varkappa_{3}\left(\epsilon z,-\frac{1}{2}-\frac{3}{2}\epsilon^{4}t_{2}-\frac{1}{2}\epsilon^{6}t_{3}\right)-3\varkappa_{2}(\epsilon z,2-2\epsilon^{4}t_{2})\right).\end{array}$
Concluding this section we note that, in the trigonometric case, sequences of the static periodic configurations (86) can be continued with the help of solutions involving the ordinary Jacobi polynomials (see (85)). In other words, in difference from the rational limit, trigonometric sequences are not terminating. However, similarly to the rational case, these sequences depend only on finite numbers of free continuous parameters.
@@ 12 Time-Dependent Darboux Transformations, Calogero-Moser Systems and Locus Configurations
As we saw in section 3, all possible static configurations of finite number of vortices of strengths ${\cal Q}_{i}=\pm 1$ are given by roots of the Adler-Moser polynomials. In the present section we will demonstrate that sequences of soliton $\tau$-functions (78,79) define all possible vortex street patterns with ${\cal Q}_{i}=\pm 1$. To do this, we will introduce dynamical systems that are completely integrable in the $\Lambda=1$ case, and whose fixed points coincide with static and translating configurations of the point vortices. These systems are related to Darboux transforms of the time-dependent free Schrodinger (or heat) equation.
@@
Here, $H$ can be presented in the form
@@
where subscript $t$ stands for time derivative. In eq. (91), $\varkappa=\varkappa(z,t)$ is one of solutions of (90). The first-order differential operators $A$, $A^{*}$ are of the same form (29) as those involved into the time-independent Darboux transformation. Permuting factors in (91), we get the new Schrodinger operator
@@
Multiplying (91) by $A$ from the left, we see that $AH=\hat{H}A-A_{t}$, and therefore $\hat{\psi}=A\psi$ solves the new equation
@@
The new potential $\hat{u}(x,t)$ and solution $\hat{\psi}$ are given by the same formulae as in the case of the time-independent Darboux transformations, i.e.
@@
Now, we take the free Schrodinger operator $H_{0}=-\partial_{z}^{2}$. The corresponding imaginary-time Schrodinger (or inverse heat) equation admits polynomial solutions $p(z,t)$
$p_{t}=-p^{\prime\prime},\quad p=\prod_{i=1}^{l}(z-z_{i}(t)),$
@@
Using $p$ as a seed function, $\varkappa=p$ in (92), we Darboux transform another polynomial solution $\psi$, $\psi_{t}=-\psi^{\prime\prime}$ into the rational solution
$\hat{\psi}=\frac{q}{p},\quad q=\psi^{\prime}p-p^{\prime}\psi.$ (93)
Then, the Schrodinger operator acquires the potential $\hat{u}=-2(\log p)^{\prime\prime}$, and the new equation $\hat{H}\hat{\psi}=\hat{\psi}_{t}$ writes, in terms of $p$ and $q$, as
$\left(-\partial_{z}^{2}-2(\log p)^{\prime\prime}\right)\left[q/p\right]=\left[q/p\right]_{t},$
@@
$qp_{t}-pq_{t}=q^{\prime\prime}p-2p^{\prime}q^{\prime}+p^{\prime\prime}q.$ (94)
Thus, we incorporated dynamics to the Tkachenko equation (9). It is not difficult to see that further iterations of Darboux transformations will lead to equations of the same form for consecutive transforms of polynomials.
Equation (94) results in the following dynamical system for roots of $p$ and $q$, denoted as $z_{1},\dots,z_{l}$ and $z_{l+1},\dots,z_{l+m}$ respectively:
@@
where negative and positive charges of equal magnitude ${\cal Q}_{i}=\pm 1$ are attached to roots of $q$ and $p$ correspondingly. If instead of polynomials $p$ and $\psi$ in (93) we take $pe^{kz-k^{2}t}$ and $\psi e^{kz-k^{2}t}$, we will get the Galilean transform of (95)
$\frac{dz_{i}}{dt}=2k+2\sum_{j=1,j\not=i}^{N}\frac{{\cal Q}_{j}}{z_{i}-z_{j}}.$ (96)
Although this system seems to resemble that of moving vortices (4), in fact, it is quite different: In contrast to (4), the system (96) contains $dz_{i}/dt$ rather than complex conjugates $d\bar{z}_{i}/dt$. As a consequence, a systems of $N$ vortices is a hamiltonian system of dimension $2N$, while system of $N$ particles (96) has dimension $N$ and is not hamiltonian. Eq. (96) has several (quite artificial) physical interpretations. For example, we can interpret it as system of sinks/sources moving with the flow of the 2d ideal fluid. This and other interpretations make sense only when all $z_{i}$ move along the same line (e.g. all $z_{i}$ are real), except cases when all particles move with equal velocities, i.e. at fixed points of (96). These fixed points coincide with translating and static configurations (5,6) of systems of vortices.
Dynamics can be also incorporated into the bi-linear equation (8) for arbitrary $\Lambda$ as
$qp_{t}-\Lambda pq_{t}=q^{\prime\prime}p-2\Lambda p^{\prime}q^{\prime}+\Lambda^{2}p^{\prime\prime}q.$
The corresponding system has the form (95), where now ${\cal Q}_{i}=-1$ for roots of $p$, and ${\cal Q}_{i}=\Lambda$ for roots of $q$.
System (95), with arbitrary ${\cal Q}_{i}$, can be embedded into a hamiltonian system of newtonian particles interacting pairwise through inverse square potentials: Taking time derivative of eq. (95), then eliminating first order time derivatives by substituting them from (95), we get
@@
(for details see e.g. [25] or [35] and references therein). For $\Lambda=1$, the sum of two charges vanishes ${\cal Q}_{i}+{\cal Q}_{j}=0$, when $i$ and $j$ belong to different species, and (97) decouples into two non-interacting sub-systems
@@
There is another case of decoupling, this time for three species of symmetrically situated particles: One particle of strength $a$ at the origin $z=0$, $2l$ particles of strength $-1$ at ±$\pm z_{1},\dots,\pm z_{l}$ and $2m$ particles of strength $1$ at ±$\pm z_{l+1},\dots,\pm z_{l+m}$. Here, due to the symmetry, we have reduction from $12(l+m)+$ 1 to $l+m$ degrees of freedom and decoupling into the two BC-type Calogero-Moser systems
@@
of $l$ and $m$ particles respectively (for review on Calogero-Moser systems related to different root systems see e.g. [34]). This case corresponds to the Darboux-transformed imaginary-time Schrodinger equation with
@@
Initial equation $H_{0}\psi=\psi_{t}$ has quasi-polynomial solutions $\psi=z^{a}\prod_{i=1}^{l}(z^{2}-z_{i}(t)^{2})$, and the Darboux transforms constructed from these solutions are quasi-rational. When $a$ is half-integer, fixed points of system (95) are configurations of even bi-spectral family from section 6.
Let us now return to the main topic of this section and consider equilibrium configurations of vortices. For these configurations $d^{2}z_{i}/dt^{2}=0$. Then form (97) it follows that the equilibrium conditions (6) imply the following $N$ locus conditions
@@
i.e. the locus conditions are necessary conditions of equilibrium. In the case of two species with ${\cal Q}_{i}=\pm 1$, locus conditions decouple, and we have the separate set of locus conditions [^16] It is worth mentioning, that the theory of algebraically integrable systems also considers loci with multiple roots, where more general locus conditions are imposed [13, 12]. for each specie

### v3

@@ 13 Equilibria of Vortex Streets - Even Family
In the previous section, we discussed vortex street configurations that are periodic generalizations of those related to the odd family of bi-spectral Schrödinger operators. Below, we introduce configurations that are periodic generalizations of those related to the even family (see section 6).
We first consider a chain of Darboux transformations that begins with $H_{0}=-\partial_{z}^{2}+u_{0}$, where $u_{0}$ is the trigonometric Poschl-Teller potential
$u_{0}=\frac{a(a-1)}{\sin^{2}z}+\frac{b(b-1)}{\cos^{2}z},$ (105)
and $a$ and $b$ are arbitrary real numbers. The simplest configuration corresponds to the eigenfunction $\sin^{a}z\cos^{b}z$ and consists of two static vortices of with strengths ${\cal Q}_{1}=a$, ${\cal Q}_{2}=b$ situated at diametrically opposite points of the $\pi$-periodic cylinder, i.e., at $z=z_{1}=0$ and $z=z_{2}=\pi/2$.
When $a$ and $b$ are integers, the potential (105) is a soliton solution of the KdV hierarchy at specific “times” (the corresponding parameters $k_{i}$ and $\zeta_{i}$ in (101, 100) can be found, e.g., in refs [11] or [23]) [^25] In general, a sequence of $n$ Darboux transformations from the chain (35), where $A_{i}=\partial_{z}-(a+i)\cot z+(b+i)\tan z$, $i=0,1,\dots$ , results in a shift of parameters $a\to a+n$, $b\to b+n$ in the potential (105).. In this case, $u_{0}$ and sequences of its Darboux transforms belong to the odd family considered in the previous section.
Due to invariance of the potential (105) under the involutions $a\to 1-a$, $b\to 1-b$, for generic $a$ and $b$ (i.e., $a\pm b\not\in\mathbb{Z}$), the Schrodinger operator $H_{0}=-\partial_{z}^{2}+u_{0}$ has four distinct types of factorizable eigenfunctions $\varkappa_{i}(z)$, $H_{0}\varkappa_{i}=\lambda_{i}\varkappa_{i}$ of the form :
@@
where $i=0,1,2,\dots$ and ${}_{2}F_{1}$ stands for the Gauss hypergeometric function. The first parameter of the ${}_{2}F_{1}$ in (106) is a non-positive integer $-i$, and the hypergeometric factors in (106) are polynomials in $\sin^{2}z$. These factorizable eigenfunctions are trigonometric-rational functions, up to a factor of $\sin^{\pm a}z\cos^{\pm b}z$. For generic $a$ and $b$, all eigenvalues $\lambda_{i}$ are distinct, and all eigenfunctions in (106) are linearly independent.
From (106), using the Wronskian formula (101), we can construct eigenfunctions of Darboux transforms $H_{n}=-\partial_{z}^{2}+u_{0}-2(\log\tau_{n})^{\prime\prime}$ of $H_{0}$. These eigenfunctions are factorizable, and, therefore, correspond to static vortex street configurations. Note that, for generic $a$ and $b$, only one of the two linearly independent solutions of the Schrodinger equation $H_{0}\varkappa_{i}=\lambda_{i}\varkappa_{i}$ is factorizable. Therefore, in contrast to the soliton-related configurations (100,101), the configurations generated from the seed functions (106) do not depend on non-trivial free parameters. Similar to the systems considered in section 11, configurations with generic $a$ and $b$ are parametrized only by a sequence of distinct integers $k_{1}<k_{2}<\dots<k_{n}$. These correspond to $\eta_{0},\eta_{1},\dots,\eta_{n-1}$ in the chain (35), where $\eta_{i}$ belong to the set of eigenvalues given in (106) . Here, generic configurations consist of four species of vortices: two species with ${\cal Q}_{i}=\pm 1$ at $z=z_{i}\not\in\{0,\pi/2\}$, a vortex of the third species at $z=0$, and a vortex of the fourth species at $z=\pi/2$.
An even family of periodic solutions emerges when $a$ and $b$ become half-integers, such that
@@
In this case, eigenvalues $\lambda_{i}$ in (106) overlap, and for some of them “degeneration” occurs [^26] Degeneration also occurs for integer $a$ and $b$. There, number of degenerate states, as we saw in the previous section, is infinite.. Now, there are two types (degenerate and non-degenerate) of eigenfunctions $\varkappa_{i}(z)$ :
$\begin{array}[]{lll}P^{(-l,-m)}_{i}(\cos 2z,\zeta_{i})\sin^{1/2-l}z\cos^{1/2-m}z,&\lambda_{i}=(2i+1-l-m)^{2},&\frac{l+m}{2}\leq i<l+m\\ P^{(l,m)}_{i}(\cos 2z)\sin^{l+1/2}z\cos^{m+1/2}z,&\lambda_{i}=(2i+1+l+m)^{2},&i\geq 0\end{array}.$ (107)
These are both trigonometric-rational up to the common factor $(\sin z\cos z)^{1/2}$. In the first type, $P^{(-l,-m)}_{i}(x,\zeta)$ is a polynomial in $x$ that also depends on a free parameter $\zeta$. Such polynomials are called para-Jacobi Polynomials [16], [13]. Solutions of the second type involve ordinary Jacobi polynomials $P^{(l,m)}_{i}$ and are non-degenerate.
@@
When $l\not=m$ (i.e., $a\not=b$), the generic configurations consist of four species of vortices: two species with ${\cal Q}_{i}=\pm 1$ at $z=z_{i}\not\in\{0,\pi/2\}$, a vortex of the third species at $z=0$, a vortex of the fourth species at $z=\pi/2$.
When $l=m$ (i.e., $a=b$), the vortices at $z=0$ and $z=\pi/2$ have the same strengths, and the generic $l=m$ configurations consist of three species of vortices.
The simplest non-trivial example of a member of the even family corresponds to $l=m=1$. Here, $a=b=3/2$, and $P^{(-1,-1)}_{1}(\cos 2z,\zeta_{1})=\cos 2z+1+\zeta_{1}$. From (107), we obtain
@@
Note that by the scaling $z\to\epsilon z$, $\psi_{1}\to-\psi_{1}/(2\epsilon^{3/2})$, and by the change of parameter $\zeta_{1}=-2(1+\epsilon^{2}s_{1})$, in the $\epsilon\to 0$ limit, we obtain
$\psi_{1}\to\lim_{\epsilon\to 0}\frac{-1}{2\epsilon^{3/2}}\sqrt{2}\frac{\cos(2\epsilon z)+1-2(1+\epsilon^{2}s_{1})}{\sqrt{\sin(2\epsilon z)}}=\frac{z^{2}+s_{1}^{2}}{z^{1/2}}.$
This is $\psi_{1}$ (see (60)) from the even bi-spectral family of section 6.
@@
$\varkappa_{m+i}=(\sin z\cos z)^{1/2-m}P^{(-m,-m)}_{m+i}(\cos 2z,\zeta_{m+i}),\quad i=0,1,\dots,m-1.$ (109)
@@
When we re-scale $z\to\epsilon z$, $H\to\epsilon^{2}H$, the eigenvalues scale as $\lambda_{i}\to\epsilon^{2}\lambda_{i}$, and in the rational limit, we have confluent eigenvalues, all tending to zero: $\lambda_{i}\to\eta=0$. According to Section 4 (see eqs. (43), (44)), in this limit, the Wronskian representation of the Darboux transforms becomes
@@
Here, $H_{0}=-\partial_{z}^{2}+\frac{m^{2}-1/4}{z^{2}}$ is the rational limit of the trigonometric Schrodinger operator (105), with $a=b=m+1/2$, obtained by the scaling $z\to\epsilon z$, $H_{0}\to\epsilon^{2}H_{0}$, i.e. $H_{0}\to\lim_{\epsilon=0}\epsilon^{2}\left(-\partial^{2}/\partial(\epsilon z)^{2}+u_{0}(\epsilon z)\right)$. In other words, ${\cal X}_{i}(z)$ are Laurent polynomials in $z^{1/2}$ that solve the following chain of equations:
${\cal X}_{1}=z^{1/2-m}(z^{2m}+s_{m}),\quad-{\cal X}_{i+1}^{\prime\prime}+\frac{m^{2}-1/4}{z^{2}}{\cal X}_{i+1}=c_{i}{\cal X}_{i},\quad i=1,2,\dots,m-1,$ (111)
where $c_{i}$ are arbitrary constants. Although two constants of integration appear at each step of the chain, only one of them is essential. The chain terminates at the $(m-1)$th step, since the logarithmic term appears at the $m$th step. The function ${\cal X}_{i}$ is the $\epsilon\to 0$ limit of a linear combinations of $i$ first eigenfunctions (109):
@@
$1,\quad\epsilon^{2m}s_{m},\quad\epsilon^{2m+2}s_{m+1},\quad\epsilon^{2m+4}s_{m+1},\dots\epsilon^{2(m+j-1)}s_{m+j}.$
The coefficients of the linear combinations are constants determined by the condition of the existence of the limit.
For example, in the case $m=2$, corresponding to the terminating sequence (61), we have
$\begin{array}[]{l}{\cal X}_{1}=\lim_{\epsilon\to 0}\epsilon^{-5/2}\varkappa_{2}(\epsilon z,2-2\epsilon^{4}s_{2}),\\ {\cal X}_{2}=\lim_{\epsilon\to 0}\epsilon^{-9/2}\left(\varkappa_{3}\left(\epsilon z,-\frac{1}{2}-\frac{3}{2}\epsilon^{4}s_{2}-\frac{1}{2}\epsilon^{6}s_{3}\right)-3\varkappa_{2}(\epsilon z,2-2\epsilon^{4}s_{2})\right).\end{array}$
Concluding this section we note that, in the trigonometric case, sequences of the static periodic configurations (108) can be continued with the help of solutions involving the ordinary Jacobi polynomials (see (107)). In other words, unlike the rational limit, trigonometric sequences are not terminating. However, similar to the rational case, these sequences depend only on a finite number of free continuous parameters.
@@ 14 Time-Dependent Darboux Transformations, Calogero-Moser Systems and Locus Configurations
As we saw in Section 3, all possible static configurations of a finite number of vortices of strengths ${\cal Q}_{i}=\pm 1$ are given by roots of the Adler-Moser polynomials. In this section we will demonstrate that sequences of soliton $\tau$-functions (100,101) define all possible vortex street patterns with ${\cal Q}_{i}=\pm 1$. To do this, we will introduce dynamical systems that are completely integrable in the $\Lambda=1$ case, and whose fixed points coincide with static and translating configurations of the point vortices. These systems are related to Darboux transforms of the time-dependent free Schrodinger (or heat) equation.
@@
Here, $H$ can be expressed in the form:
@@
where subscript $t$ denotes the time derivative. In equation (113), $\varkappa=\varkappa(z,t)$ is one of solutions of (112). The first-order differential operators $A$ and $A^{*}$ are of the same form as in equation (30). Permuting factors in (113), we obtain the new Schrodinger operator
@@
Multiplying (113) by $A$ from the left, we obtain $AH=\hat{H}A-A_{t}$, and therefore $\hat{\psi}=A\psi$ solves the new equation
@@
The new potential $\hat{u}(x,t)$ and solution $\hat{\psi}$ are given by the same formulae as in the case of the time-independent Darboux transformations, i.e.,
@@
Now, we take the free Schrodinger operator $H_{0}=-\partial_{z}^{2}$. The corresponding imaginary-time Schrodinger (or inverse heat) equation admits polynomial solutions $q(z,t)$:
$q_{t}=-q^{\prime\prime},\quad q=\prod_{i=1}^{l}(z-z_{i}(t)),$
@@
Using $q$ as a seed function ($\varkappa=q$ in Eq. (114)) we can apply the Darboux transformation to another polynomial solution $\psi$, $\psi_{t}=-\psi^{\prime\prime}$. The transform $\hat{\psi}$ of $\psi$ is a rational function:
$\hat{\psi}=\frac{p}{q},\quad p=\psi^{\prime}q-q^{\prime}\psi.$ (115)
The Schrodinger operator acquires the potential $\hat{u}=-2(\log p)^{\prime\prime}$, and the new equation $\hat{H}\hat{\psi}=\hat{\psi}_{t}$ writes, in terms of $p$ and $q$, as
$\left(-\partial_{z}^{2}-2(\log q)^{\prime\prime}\right)\left[p/q\right]=\left[p/q\right]_{t},$
@@
$pq_{t}-qp_{t}=p^{\prime\prime}q-2p^{\prime}q^{\prime}+q^{\prime\prime}p.$
Thus, we have incorporated dynamics into the Tkachenko equation (11). It is easy to see that further iterations of Darboux transformations will lead to equations of the same form for consecutive transforms of polynomials.
The time-dependent Tkachenko equation results in the following dynamical system for the roots of $p$ and $q$, denoted as $z_{1},\dots,z_{l}$ and $z_{l+1},\dots,z_{l+m}$ respectively:
@@
where negative and positive charges of equal magnitude ${\cal Q}_{i}=\pm 1$ are attached to roots of $p$ and $q$, respectively.
System (116), with arbitrary [^28] This is a system for the roots of polynomials ${\cal P}_{i}(z,t)$, which satisfy the time-dependent generalization of the poly-linear equation (71): $-\sum_{i}\Lambda_{i}\frac{\partial{\cal P}_{i}/\partial t}{{\cal P}_{i}}=\sum_{i}\Lambda_{i}^{2}\frac{{\cal P}_{i}^{\prime\prime}}{{\cal P}_{i}}+2\sum_{i<j}\Lambda_{i}\Lambda_{j}\frac{{\cal P}_{i}^{\prime}}{{\cal P}_{i}}\frac{{\cal P}_{j}^{\prime}}{{\cal P}_{j}}$, see [32]. ${\cal Q}_{i}$, can be embedded into a hamiltonian system of newtonian particles interacting pairwise through inverse square potentials: By taking the time derivative of eq. (116), and then eliminating the first order time derivatives by substituting them from (116), we obtain:
@@
(For details, see e.g. [32] or [44]). For $\Lambda=1$, the sum of two charges vanishes, ${\cal Q}_{i}+{\cal Q}_{j}=0$, when $i$ and $j$ belong to different species, and (117) decouples into two non-interacting sub-systems
@@
There is another case of decoupling, this time for three species of symmetrically situated particles: One particle of strength $a$ at the origin $z=0$, $2l$ particles of strength $-1$ at ±$\pm z_{1},\dots,\pm z_{l}$ and $2m$ particles of strength $1$ at ±$\pm z_{l+1},\dots,\pm z_{l+m}$. Here, due to the symmetry, the system reduces from $12(l+m)+$ 1 to $l+m$ degrees of freedom and decoupling into the two BC-type Calogero-Moser subsystems
@@
of $l$ and $m$ particles respectively [^29] For a review of Calogero-Moser systems related to different root systems, see e.g. [43].. This case corresponds to the Darboux-transformed imaginary-time Schrodinger equation with
@@
The initial equation $H_{0}\psi=\psi_{t}$ has quasi-polynomial solutions of the form $\psi=z^{a}\prod_{i=1}^{l}(z^{2}-z_{i}(t)^{2})$, and the Darboux transforms constructed from these solutions are quasi-rational. When $a$ is a half-integer, the fixed points of system (116) correspond to configurations of the even bi-spectral family from Section 6.
Let us now return to the main topic of this section and consider equilibrium configurations of vortices. For these configurations $d^{2}z_{i}/dt^{2}=0$. Then form (117) it follows that the equilibrium conditions (6) imply the following $N$ locus conditions [^30] Note that additional locus conditions for the case ${\cal Q}_{i}\in\{-1,2\}$ were found by O’Neil [41].
@@
i.e., the locus conditions are necessary conditions for equilibrium. In the case of two species with ${\cal Q}_{i}=\pm 1$, locus conditions decouple, resulting in a separate set of locus conditions [^31] It is worth mentioning that the theory of algebraically integrable systems also considers loci with multiple roots, where more general locus conditions are imposed [17, 18]. for each species: